OpenStudy (anonymous):

4x + 2(3x – 9) = x A. x = –2 B. x = 1.8 C. x = 2 D. x = 18

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

i think its B but im not sure

OpenStudy (anonymous):

c. x=2

OpenStudy (anonymous):

thank you @borimexz

OpenStudy (anonymous):

wait becuase when you plug in 2 for x it equals to -2

OpenStudy (anonymous):

so A @borimexz

OpenStudy (anonymous):

4x+2(3x-9)=x 4x+6x-18=x +18 +18 10x=x+18 -x -x 9x=18 /9 /9 x=2

OpenStudy (anonymous):

alright thank you :) @borimexz